Is this answer correct if I am trying to find the value of "t" in this equation? Here is the problem

{ { { t/8 - 6 = -12 } } }

and I got " t = 48 "

Answer by rfer(16322)

t/8-6=-12
t/8=-12+6
t/8=-6
t=-6*8
t=-48
Watch your signs